OpenStudy (anonymous):

Given f(x) and g(t). Prove that if f(x)=g(t) then that f(x)=a and g(t)=a where a is the same constant.

OpenStudy (anonymous):

assume that f(x) = a for some value of x. since f(x) =g(t) then a=g(t) so f(x) and G9t)=a

OpenStudy (anonymous):

Lets say that we didn't know there was a constant. Assume we only know that f(x)=g(t)

OpenStudy (anonymous):

assume that f(x) = a for some value of x. since f(x) =g(t) then a=g(t) so f(x) and G9t)=a

OpenStudy (anonymous):

I am not sure that I follow. The question states that a is a constant.

OpenStudy (anonymous):

@thomasj Well I was solving this problem where I end up with two functions being equal to each other. Then my book states that because the functions are equal to each other that the functions are equal to a constant.

OpenStudy (anonymous):

why would it be constant? let \(f(x)=x^2=g(x)\) then f and g are not constant. maybe there was something else in the problem?

OpenStudy (anonymous):

@satellite73 No the functions are not of the same variable. one is of x and the other one is of t

OpenStudy (anonymous):

variable is not important in a function. there is no difference between \(f(x)=x^2\) and \(f(z)=z^2\) it just means square right?

OpenStudy (anonymous):

I'm it does matter in this case. x refers to distance where t refers to time.

OpenStudy (anonymous):

Or does that not matter at all I'm not sure.